OpenStudy (anonymous):

an you please show me how to solve an equation like this?! It's not making sense. x² + 6x = –5

OpenStudy (abb0t):

Add the 5 and set your polynomial to zero.

OpenStudy (anonymous):

x² + 6x + 5 = 0?

OpenStudy (abb0t):

\[x^2+6x+5 =0\] notice that it's in the form: \[ax^2+bx+c\] to factor, you must find TWO numbers whose product is "c", but the sum of the two terms, result in "b". Does that make sense? It means: n x n = c, but n + n = b

OpenStudy (abb0t):

where the n's are two constants. meaning numbers - obviously.

OpenStudy (anonymous):

Um...still don't get it.

OpenStudy (anonymous):

you have to find two numbers whose product is 5 and sum is 6

OpenStudy (anonymous):

Oh! Okay!

OpenStudy (abb0t):

Yes.

OpenStudy (anonymous):

But...that's for factoring right? I need to find the value of x...

OpenStudy (anonymous):

yea, you can solve for x by factoring

OpenStudy (abb0t):

Yes, to find the value(s) of "x". You need to factor.

OpenStudy (anonymous):

or you can use the quadratics formula

OpenStudy (anonymous):

Okay...

OpenStudy (abb0t):

quadratic formula usually makes the algebra more tedious. It's best if you factor.

OpenStudy (anonymous):

Can you guys check my answer once I attempt to figure it out?

OpenStudy (anonymous):

sure

OpenStudy (abb0t):

Of course.

OpenStudy (anonymous):

Okay, I'm coming up with jibberish. Could you walk me through it?

OpenStudy (anonymous):

make a list of factors of 5

OpenStudy (anonymous):

just 1 and 5

OpenStudy (anonymous):

and -1 and -5

OpenStudy (anonymous):

yes. so (x + 1)(x + 5) = 0 now you can set those two factors (x + 1 and x + 5) set to 0 and solve for x

OpenStudy (anonymous):

right, but you couldn't use -1 and -5 because those don't add up to 6

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so solve for x in both x + 5 = 0 and x + 1 = 0 and they will be your answers

OpenStudy (anonymous):

-5 and -1?

OpenStudy (anonymous):

yes :]

OpenStudy (anonymous):

okay thanks!

OpenStudy (anonymous):

no problem